Required information For a bolted assembly with six bolts, the stiffness of each bolt is kb = 2.6 Mlbf/in and the stiffness oI need help Please, third time posting this question and i keep getting the wrong answer.
Answer #1

Feel free to ask any question but remember to rate the answer.

Giyen - P= 80 kips Kb= 2.6 Mebflin Km=is Mebflin Drameter of bolts = Lin -13 UNC grade 8 No. of bolts= 6 for ļ in-13 UNC grad

Preload stress, 61=0.75 Sp = 0.75X120 go kpsi Stiffness constant, C= kb Kot km C= 2.6 (2.6+15) C=0.1477 Yielding factor of sa